OpenStudy (anonymous):

Show that as x->0, sinx = x - x^3/3! + o(x^4)

OpenStudy (michele_laino):

that is the Taylor expansion of the function sin x

OpenStudy (anonymous):

what is taylor expansion?

OpenStudy (michele_laino):

It is a power series, here are the first three terms of the Taylor expansion: \[\Large \begin{gathered} \sin x = \sin \left( 0 \right) + {\left. {\frac{{d\sin x}}{{dx}}} \right|_{x = 0}}x + \frac{1}{{2!}}{\left. {\frac{{{d^2}\sin x}}{{d{x^2}}}} \right|_{x = 0}}{x^2} + \hfill \\ \hfill \\ + {\left. {\frac{1}{{3!}}\frac{{{d^3}\sin x}}{{d{x^3}}}} \right|_{x = 0}}{x^3} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

for example, we have: \[\Large {\left. {\frac{{d\sin x}}{{dx}}} \right|_{x = 0}} = \cos \left( 0 \right) = 1\]

OpenStudy (anonymous):

so would it be, sinx = x- x^3/3! and then because limit x->0 (sinx-x-x^3/3!) = lim x->( sinx-sinx) = 0, then sinx = x-x^3/x! + o(x^4)?